OpenStudy (wondermath):

What is the 1000th derivative of y=e^x(sinx)?

OpenStudy (thomas5267):

Which of the following is the question? A or B? \[ \begin{align*} \text{A.}\quad &y=e^{x(\sin(x))}\\ \text{B.}\quad &y=\sin(x)e^x \end{align*} \]

OpenStudy (wondermath):

its B, sorry

OpenStudy (wondermath):

i think i got the pattern going

OpenStudy (wondermath):

let me type out what i have

OpenStudy (wondermath):

y'= (e^x)(sinx+cosx) y''=2(e^x)(cosx) y'''=2(e^x)(cosx-sinx) y''''= (-2^2)(e^x)(sinx)

OpenStudy (anonymous):

\[y^{1000} =2y ^{999} -2y ^{998}\]

OpenStudy (anonymous):

\[(-1)^{250} 2^{500} e^x \sin (x) \] In general if k=4m, the kth derivative is \[ (-1)^m 2^{2 m} e^x \sin (x) \]

OpenStudy (thomas5267):

Astonishing! @eliassaab

OpenStudy (wondermath):

the answer is (2^500)(e^x)(sinx), but idk how they got it

OpenStudy (anonymous):